OpenStudy (anonymous):

line is passing throught (8,16) and (9,18) Find point slope form and point interscept form

OpenStudy (anonymous):

m=(y1-y)/(x1-x) m=(18-16)/(9-8) what is the slope m =?

OpenStudy (anonymous):

they have not given m

OpenStudy (anonymous):

if m=2 then equation of the line: y-16=2(x-8) y=2x

OpenStudy (anonymous):

not given... you calculate it - see above

OpenStudy (anonymous):

what is the point slope form

OpenStudy (anonymous):

The other format for straight-line equations is called the "point-slope" form. For this one, they give you a point (x, y) and a slope m, and have you plug it into this formula: y1-y=m(x1-x)

OpenStudy (anonymous):

check this site: http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (anonymous):

your answer: y=2x (see solution above)

OpenStudy (anonymous):

so what would be the slope intercept form

OpenStudy (anonymous):

y=2x

OpenStudy (anonymous):

i mean slope intercept form